# STANCE-$\text{C}^{3}$: Domain-adaptive Cross-target Stance Detection via Contrastive Learning and Counterfactual Generation

### v1

@@ Abstract
Stance detection is the process of inferring a person’s position or standpoint on a specific issue to deduce prevailing perceptions toward topics of general or controversial interest, such as health policies during the COVID-19 pandemic. Existing models for stance detection are trained to perform well for a single domain (e.g., COVID-19) and a specific target topic (e.g., masking protocols), but are generally ineffectual in other domains or targets due to distributional shifts in the data. However, constructing high-performing, domain-specific stance detection models requires an extensive corpus of labeled data relevant to the targeted domain, yet such datasets are not readily available. This poses a challenge as the process of annotating data is costly and time-consuming. To address these challenges, we introduce a novel stance detection model coined domain-adaptive Cross-target STANCE detection via Contrastive learning and Counterfactual generation (STANCE-$\text{C}^{3}$) that uses counterfactual data augmentation to enhance domain-adaptive training by enriching the target domain dataset during the training process and requiring significantly less information from the new domain. We also propose a modified self-supervised contrastive learning as a component of STANCE-$\text{C}^{3}$ to prevent overfitting for the existing domain and target and enable cross-target stance detection. Through experiments on various datasets, we show that STANCE-$\text{C}^{3}$ shows performance improvement over existing state-of-the-art methods.
@@ I Introduction
Inferring prevailing public attitudes toward specific personalities, policies, or events, especially those with controversial undertones, is one of the many applications of mining social media data. In natural language processing (NLP), the task of predicting the general public’s outlook for specific target keyword(s) is called stance detection. The target represents the subject or topic of interest, while the stance reflects the prevailing attitude or position. As an example from the tweet “Everyone should get the COVID-19 vaccine”, a positive stance is posited toward COVID-19 vaccine as the target. Stance detection has had diverse applications in tasks such as investigating users’ opinions on a new product [1], understanding public acceptance of a newly proposed legislation [2], and determining support for public health measures [3].
Stance detection models have been proposed under two constructs, namely, the single-target and cross-target scenarios. The single-target scenario determines users’ opinions toward a single, fixed target [4], while cross-target methods extend the model’s versatility to address multiple targets [5]. In single-target cases, the training set includes labeled data related to the target, while in cross-target methods, there is lacking or very few labeled data on the target of interest.
In addition, current state-of-the-art (SOTA) stance detection models are primarily focused on approaches for improving the model’s performance on a single domain. Previous research has illustrated, however, that these models do not necessarily perform well when used in a domain that is distinct from the domain of the training data set  [6]. We hypothesize that this performance degradation is mainly due to the following attributes of current SOTA stance detection models: (1) the tendency for attention bias toward events related to the given target, a phenomenon that we will illustrate in experiments in later sections; (2) the absence of target words in predicting stance during the training phase; and (3) the model’s reliance on auxiliary words (i.e., words that are not necessarily the target or stance) specific to a given domain. Additionally, the scarcity of available training data limits the capabilities of stance detection models across multiple domains and targets.
Table I shows an example of a domain-adaptive cross-target stance detection problem. In this example, the two domains differ by hashtags and data acquisition time periods, while the targets are distinguished by the topical content (vaccination vs. mask wearing). Other attributes of the sampling process could account for distinctions among domains and targets, such as data sources (e.g., social media vs. news) or content genre (political vs. gossip news).
TABLE I: Examples of domain-adaptive cross-target stance detection
Source Target:
COVID-19 vaccination
Statement: Got my first vaccine dose. I’m so thankful for all the healthcare workers that made the experience positive.
Stance: In favor of
Source Domain:
Tweets with target-related keywords (e.g. get vaccinated) from January 1st, 2020 to August 23rd, 2021
Destination Target:
wearing face masks
Statement: I will not take the mask and I will not wear a mask. In the words of Patrick Henry "Give me liberty or give me death!"
Stance: Against
Target Domain:
Tweets with target-related hashtags (e.g. #MasksSaveLives) from February 27th, 2020 to August 20th, 2020
@@
In this paper, we propose a cross-domain, cross-target stance detection model called domain-adaptive Cross-target STANCE detection via Contrastive learning and Counterfactual generation (STANCE-$\text{C}^{3}$) to address limitations in existing SOTA models. STANCE-$\text{C}^{3}$ is a domain- and target-agnostic model that leverages counterfactual data augmentation (CDA) [7], an approach that has been shown to decrease bias and improve the robustness of neural network-based models [8]. We combine CDA with text style transfer methods [9] to facilitate effective text transfer from one domain to another. These language model-based data augmentation techniques are expected to enrich existing datasets with information that will improve the performance of stance detection models across multiple domains.
@@
Previous work on stance detection addressed domain robustness by using adversarial training and including the target word as an input to the stance detection model (e.g., concatenating the target word and text as input to the language model). However, this approach resulted in overfitting the model to the specific target, thus limiting its performance robustness for other targets. In STANCE-$\text{C}^{3}$, we adopt a self-supervised technique called contrastive learning [10] to extract both shared and distinct high-level features that capture the specific characteristics of statement-target pairs. Text features such as word embeddings, Part-of-Speech tags, or $n$-grams are commonly used across tasks or datasets, while domain-specific terms, phrases, or concepts are normally used for target topics or subjects. In addition to the existing contrastive loss function, we incorporate additional components such as cosine similarity and negative pair loss to guide the model to minimize the distance between pairs with same stance labels and maximize the distance for pairs with different stance labels. Because contrastive learning is classified as a self-supervised approach, STANCE-$\text{C}^{3}$ benefits from requiring fewer manually annotated data points. Additionally, when combined with CDA during the pre-processing stage, it ensures that the training set is optimally constructed for a cross-domain and cross-target setting, i.e., there are sufficient data points exhibiting both similarities and differences, allowing the model to infer salient features that are useful for stance detection.
The major contributions of this work are as follows: (1) we investigate the problem of domain-adaptive cross-target stance detection and propose a novel solution for training a model when data on the target domain and target object (or destination target) are sparse; (2) we demonstrate the utility of contrastive learning as a possible solution for improving a model’s robustness across domains; and finally, (3) we provide empirical evidence of STANCE-$\text{C}^{3}$’s effectiveness by running experiments and ablation studies on real-world datasets.
II Problem Statement
@@
Suppose we have two sets of instances $\mathcal{D}_{\mathcal{S}}$ and $\mathcal{D}_{\mathcal{T}}$ from source domain $\mathcal{S}$ and target domain $\mathcal{T}$, respectively. The set $\mathcal{D}_{\mathcal{S}}={(r_{\mathcal{S}},t^{o}_{,}y_{\mathcal{S}})}^{N}$ contains N annotated instances from a source domain $\mathcal{S}$ where each instance $r_{\mathcal{S}}$ is a text (e.g., tweet) consisting of a sequence of k words, and $y_{\mathcal{S}}$ is the stance toward a target $t^{o}$ associated with each annotated instance. Another set $\mathcal{D}_{\mathcal{T}}=(r_{\mathcal{T}},t^{d})^{N^{\prime}}$ is a set of N’ instances from target domain $\mathcal{T}$ with the destination target $t^{d}$. The goal of the domain-adaptive cross-target stance detection is to learn the stance classifier $F$ that uses the features of the source domain $\mathcal{S}$ and target words $t^{o}$ in $\mathcal{D}_{\mathcal{S}}$ and predict the stance label $y_{\mathcal{T}}$ of each text $r_{\mathcal{T}}$ in $\mathcal{D}_{\mathcal{T}}$. We formally define the problem as follows:
@@
Definition (Domain Adaptive Cross-Target Stance Detection). Given statements from two separate datasets $\mathcal{D}_{\mathcal{S}}$ and $\mathcal{D}_{\mathcal{T}}$, and corresponding targets $t^{o}$ and $t^{d}$ from the source $\mathcal{S}$ and target $\mathcal{T}$ domains, respectively, learn a domain and target-agnostic text representation using $\mathcal{D}_{\mathcal{S}}$ and a small portion of labeled $\mathcal{D}_{\mathcal{T}}$ that can be classified correctly by the stance classifier $F$.
III Proposed Model
Figure 1: The proposed STANCE-$\text{C}^{3}$ architecture consists of two main components: counterfactual data generation, and contrastive learning networks. First, the counterfactual data generation network aims to enhance generalization by augmenting the dataset. In this process, tokens are represented as numeric vectors and a T5-based language model is trained to generate examples with the same sentence structure but different semantic context (i.e. domain). These generated examples are referred to as domain counterfactuals. Then using the augmented data, the contrastive learning network learns a cross-target representation that can be used for domain-adaptive cross-target stance detection.
@@ III Proposed Model
In this section, we describe our proposed model STANCE-$\text{C}^{3}$. As shown in Figure 1, the architecture consists of two main components: (1) a domain counterfactual generator that is trained with input representations of the labeled samples from the source and a small portion from the target domain and generates a set of domain counterfactuals. The generated counterfactuals provide knowledge of the target domain during the stance classification training phase. (2) These samples are then combined with the original dataset (i.e., source domain dataset with a small portion of target domain dataset) to train a BERT-based stance classifier using contrastive learning to create a domain-adaptive cross-target stance detection classifier. We split a small number of randomly selected samples from the combined datasets into positive and negative pairs according to their stance regardless of the target. This approach for contrastive loss enables us to improve the stance detection method’s generalization to unseen targets. In the following subsections, we explain the two stages in detail.
III-A Domain-adaptive Single-target
@@
One major challenge in domain-adaptive stance detection is that collecting samples for a target domain can be expensive and usually requires human annotations. Even with annotated data, modeling a domain-adaptive stance detection is difficult due to the complexity of the dataset which contains far more distinct features over different domains [11]. These features often form spurious correlations within a specific domain and make the models brittle to domain shift. For example, the presence of question marks in a specific stance or highly negative sentiment associated with a particular stance leads to a decrease in performance [6].
@@
To address the aforementioned challenges and control these variants on the domain-adaptive stance detection problem, we adapt counterfactual concepts to the source domain. A domain-counterfactual example is defined as a result text of intervening on the domain variable of the original example and converting it to the target domain while keeping other features (e.g., overall structure of the sentence) fixed. Given an example data from the COVID-19 vaccination domain (source domain), the generator first recognizes the terms related to the domain. Then, it intervenes on these terms, replacing them with text that links the example to the wearing face mask domain (target domain) while keeping the stance. We utilize source domain instances in combination with a small portion of target domain instances to incorporate target domain information in generating the counterfactuals specific to the target domain. These counterfactuals are then combined as input for the subsequent stage. Following the work of  [12], we adopt a T5-based language model (LM) to generate counterfactuals for a given target domain $\mathcal{T}$. In this approach, we train the T5-based LM to convert samples from source domain $\mathcal{S}$ to a given target domain $\mathcal{T}$.
@@
Generating domain counterfactuals consists of two main steps - domain corruption and reconstruction. In the first step of domain corruption, we get the masking score of all n-grams $w$ ($n\in\{1,2,3\}$) in a dataset $\mathcal{D}_{\mathcal{S}}$ following [12]. The affinity of $w$ to domain $\mathcal{S}$ is defined as $\rho(w,\mathcal{S})=P(\mathcal{S}|w)(1-\frac{H(\mathcal{S}|w)}{\log N}),$ where $H(\mathcal{S}|w)$ is the entropy of $\mathcal{S}$ given each n-gram $w$ and $N$ is the number of unlabeled domains we know. The final masking score given source domain $\mathcal{S}$ and target domain $\mathcal{T}$ is $mask(w,\mathcal{S},\mathcal{T})=\rho(w,\mathcal{S})-\rho(w,\mathcal{T})$. Here, a higher score implies the n-gram is highly related to the source domain and distant from the target domain.
@@
The next step is to reconstruct the masked examples $M(x)$ by concatenating them with a trainable embedding matrix, domain orientation vectors. The matrix is initialized with the embedding vector of each domain name and domain representative words. The concatenated matrix is trained with an encoder-decoder T5 architecture. Note that $\mathcal{T}=\mathcal{S}$ and $mask(w,\mathcal{S},\mathcal{T})=0$ during the training process. Given the target domain and its orientation vectors in the test phase, the trained model generates domain counterfactual $x^{\prime}$.
III-B Domain-adaptive Cross-target

### v2

# Robust Stance Detection: Understanding Public Perceptions in Social Media

@@ Abstract
The abundance of social media data has presented opportunities for accurately determining public and group-specific stances around policy proposals or controversial topics. In contrast with sentiment analysis which focuses on identifying prevailing emotions, stance detection identifies precise positions (i.e., supportive, opposing, neutral) relative to a well-defined topic, such as perceptions toward specific global health interventions during the COVID-19 pandemic. Traditional stance detection models, while effective within their specific domain (e.g., attitudes towards masking protocols during COVID-19), often lag in performance when applied to new domains and topics due to changes in data distribution. This limitation is compounded by the scarcity of domain-specific, labeled datasets, which are expensive and labor-intensive to create. A solution we present in this paper combines counterfactual data augmentation with contrastive learning to enhance the robustness of stance detection across domains and topics of interest. We evaluate the performance of current state-of-the-art stance detection models, including a prompt-optimized large language model, relative to our proposed framework succinctly called STANCE-C3 (domain-adaptive Cross-target STANCE detection via Contrastive learning and Counterfactual generation). Empirical evaluations demonstrate STANCE-C3’s consistent improvements over the baseline models with respect to accuracy across domains and varying focal topics. Despite the increasing prevalence of general-purpose models such as generative AI, specialized models such as STANCE-C3 provide utility in safety-critical domains wherein precision is highly valued, especially when a nuanced understanding of the concerns of different population segments could result in crafting more impactful public policies.
@@ I Introduction
The influence of public perceptions on policy development and societal outcomes is well-established in the context of public health [1], legislation [2], environmental sustainability [3], and other community-focused sectors. While policymakers from previous generations relied on expensive surveys or town halls to gauge public opinion toward a specific issue, the meteoric rise of social media has facilitated the swift and cost-effective collection of user-generated insights, which are voluntarily and freely provided by netizens. At the height of COVID-19, for example, social media data and natural language processing (NLP) models were leveraged to establish associations between public support for non-pharmaceutical interventions and the decreased spread of the virus [4].
NLP models that infer the collective attitude or stance toward specific targets using corpora of user-provided data from online forums (e.g., X) are called stance detection models. Unlike sentiment analysis, which detects general emotional responses toward a topic, stance detection models identify positions toward a target by analyzing the context, linguistic nuances, and even the implied meanings within the text. For instance, consider an online post stating, “I hate having to wear a mask every day, but I totally agree it’s essential to stop the spread of COVID-19." Sentiment analysis might classify this as negative due to the expressions of dislike and discomfort. However, despite the negative emotional tone, the stance towards the target (mask-wearing as a COVID-19 health measure) is clearly supportive. Consequently, stance detection offers significant opportunities to leverage social media data for accurately identifying public stances towards various initiatives. This can substantially enhance policy-making, refine strategies that impact the public, and contribute to a more informed public dialogue [5].
(a) Source Domain
(b) Target Domain
Figure 1: Examples of domain-adaptive cross-target stance detection. (a) Tweets from source domain are collected using target-related keywords/hashtags (e.g., get vaccinated) from January 1st, 2020 to August 23rd, 2021, whereas (b) tweets from target domain are collected with target-related hashtags (e.g., #MasksSaveLives) from February 27th, 2020 to August 20th, 2020.
Figure 2: The STANCE-C3 architecture consists of two key components: a counterfactual data generation network and a contrastive learning network. The counterfactual network (left) is built on a T5-based model and generates training examples that maintain sentence structure but has increased diversity in semantic context. The contrastive learning network (right) then uses the augmented dataset to learn cross-target representations by minimizing distances between examples with the same stance (positive pairs) and maximizing distances between examples with different stance (negative pairs). This approach helps acquire domain-invariant features, improving stance detection across targets.
One of the significant challenges in leveraging social media data for stance detection is the limited adaptability of existing models across different domains. This limitation is particularly acute in rapidly changing situations, where shifting public opinion across various domains (e.g., different time points) can significantly influence responses to crises, such as public health emergencies. The social cost of misinformation and the urgent need for responsive public opinion analysis across diverse domains highlight a pressing challenge: developing stance detection models that are both robust and adaptable to different contexts.
Current state-of-the-art (SOTA) stance detection models are largely focused on improving the model’s performance within a single domain. Yet, these models often underperform when applied to domains distinct from their training datasets [6]. We attribute this performance degradation to several factors: (1) Attention bias toward events related to the given target, a phenomenon we will illustrate with experiments in later sections; (2) The absence of target words in stance prediction during the training phase; and (3) The model’s reliance on auxiliary words (i.e., words that are not necessarily the target or stance) specific to a given domain. Further, the limited availability of training data restricts these models’ effectiveness across multiple domains and targets. Although Large Language Models (LLMs) have gained popularity, their resource-intensive and general-purpose nature pose issues in their usefulness for stance detection. LLMs demand substantial computational resources for training and inference, making them financially and environmentally costly. For specific tasks, smaller models can achieve comparable or even better performance without the resource burden [7, 8]. Additionally, human factors significantly influence the inference process when using LLMs; for example, it is well-established that different prompting techniques could yield widely varying results [9].
Stance detection models can also be differentiated based on the number of intended targets. Single-target models identify positions toward a specific, fixed target [10], whereas cross-target models extend the model’s versatility to multiple targets [11]. Approaches for single-target stance detection use training sets that only include labeled data related to the specific target. In contrast, cross-target methods address the challenge of the sparseness in target-specific labeled data, leveraging additional information about the new target for accurate stance predictions.
With social media serving as the main platform where people express their opinions and grievances, the development of robust stance detection methods is crucial for accurately gauging public opinion, deterring misinformation, and facilitating constructive online dialogue. Single-target models often struggle to adapt to the quickly evolving landscape of public discourse, where topics such as climate change, social justice, and public health are deeply interwoven in their influence and digital presence. Cross-target stance detection models offer a more comprehensive understanding of public opinions across related topics, eliminating the need to rebuild models for each new issue. This not only makes AI tools more responsive and adaptable but also improves their ability to provide real-time insights across a wide range of issues. Such capabilities are particularly valuable in fields like policy making and public health, where timely and informed decision-making is crucial. For example, a cross-target model could analyze opinion on related non-pharmaceutical health interventions like vaccination and mask-wearing, despite varying domains and data sources, enabling more effective responses to public health crises. Figure 1 shows an example of a domain-adaptive cross-target stance detection problem, illustrating differences in domains by search query and data acquisition time periods, while targets differ in topical content (e.g., vaccination vs. mask-wearing). Other attributes of the sampling process could account for distinctions among domains and targets, such as data sources (e.g., social media vs. news), content genre (e.g., political vs. gossip news), or publication type (e.g., opinion pieces vs. reportage).
@@
In this paper, we introduce STANCE-C3 as a domain- and target-robust model that uses Counterfactual Data Augmentation (CDA) [12], an approach that has been shown to decrease bias and improve the robustness of neural network-based models [13]. We combine CDA with text style transfer methods [14] to facilitate effective text transfer from one domain to another. These language model-based data augmentation techniques are expected to enrich existing datasets, improving the performance of stance detection models across multiple domains.
@@
Previous work on stance detection addressed domain robustness using adversarial training and incorporating the target word as an input to the model. However, this approach often led to overfitting on the specific target, restricting its robustness in performance for other targets. In STANCE-C3, we modify a self-supervised technique called contrastive learning [15] to extract both shared and distinct high-level features to capture specific characteristics of statement-target pairs. We incorporate components such as cosine similarity and negative pair loss to direct the model to minimize the distance between pairs with the same stance labels and maximize the distance between pairs with different stance labels. By requiring fewer manually annotated data points and combining CDA during the pre-processing stage, STANCE-C3 ensures that the training set is optimally constructed for a cross-domain and cross-target setting (i.e., a sufficient balance between data points that exhibit similarities and differences, allowing the model to infer salient features that are useful for stance detection) while avoiding overfitting.
The major contributions of this work are threefold: firstly, it addresses the problem of domain-adaptive cross-target stance detection, offering a novel approach to train models effectively even when data on the target domain and subject are limited. Secondly, it tests the potential of a modified contrastive loss learning method as a more effective alternative for improving model robustness in classification tasks. Lastly, it demonstrates STANCE-C3’s efficacy and improvement over existing SOTA stance detection models through rigorous experiments and ablation studies on real-world datasets.
II Related Work
STANCE-C3 is proposed in this paper to improve existing models for stance detection and robustness across domains. In this section, we review current NLP SOTA models for stance detection and domain adaptation.
II-A Stance Detection
There are two general categories of distinctions among stance detection models. Firstly, models are distinguished on how stance features are modeled. Content-level approaches use linguistic features such as topics [16] or targets [17] alongside sentiment information [18]. User-level approaches focus on user-related attributes interactions [19], information [20], and timelines [21]. Hybrid models combine content- and user-level features for post representations [22]. In this paper, we focus on content-level modeling with target embedding using BERT due to its practicality in data acquisition.
Secondly, stance detection models differ concerning the specificity of the target of interest. While many previous works focus on specific single-target scenarios [23], recent studies have considered cross-target [17], multi-target [24], and few-shot or zero-shot stance detection [25]. While both cross-target and zero-shot stance detection models are trained on one or more source targets and used on previously unseen targets [26], zero-shot methods can predict stances for unseen targets that are not necessarily related to the target in the training set [27]. Both attempt to extract transferable knowledge from sources using methods such as hierarchical contrastive learning [28]. STANCE-C3 uses a similar but modified approach to the contrastive loss during training.
II-B Domain Adaptation
As a category of transfer learning, domain Adaptation (DA) leverages agnostic or common information between domains with the target task remaining constant [29]. There are several DA setups, including unsupervised domain adaptation (UDA) where both labeled source data and unlabeled target data are available [30], semi-supervised domain adaptation (SSDA) where labeled source data and a small number of labeled target data are accessible [31], and supervised domain adaptation (SDA) where both labeled source and target data are available during training [32]. Our focus is on SSDA, where a smaller set of labeled target data is used in the training process.
III Problem Statement
@@
Suppose we have two sets of instances $\mathcal{D}_{\mathcal{S}}$ and $\mathcal{D}_{\mathcal{T}}$ from source domain $\mathcal{S}$ and target domain $\mathcal{T}$, respectively. The set $\mathcal{D}_{\mathcal{S}}={(r_{\mathcal{S}},t^{o},y_{\mathcal{S}})}^{N}$ contains N annotated instances from a source domain $\mathcal{S}$ where each instance $r_{\mathcal{S}}$ is a text (e.g., tweet) consisting of a sequence of k words, and $y_{\mathcal{S}}$ is the stance toward a target $t^{o}$ associated with each annotated instance. Another set $\mathcal{D}_{\mathcal{T}}=(r_{\mathcal{T}},t^{d})^{N^{\prime}}$ is a set of N’ instances from target domain $\mathcal{T}$ with the destination target $t^{d}$. The goal of the domain-adaptive cross-target stance detection is to learn the stance classifier $F$ that uses the features of the source domain $\mathcal{S}$ and target words $t^{o}$ in $\mathcal{D}_{\mathcal{S}}$ and predict the stance label $y_{\mathcal{T}}$ of each text $r_{\mathcal{T}}$ in $\mathcal{D}_{\mathcal{T}}$. We formally define the problem as:
@@
Definition Given statements from two separate datasets $\mathcal{D}_{\mathcal{S}}$ and $\mathcal{D}_{\mathcal{T}}$, and corresponding targets $t^{o}$ and $t^{d}$ from the source $\mathcal{S}$ and target $\mathcal{T}$ domains, respectively, learn a domain and target-agnostic text representation using $\mathcal{D}_{\mathcal{S}}$ and a small portion of labeled $\mathcal{D}_{\mathcal{T}}$ that can be classified correctly by the stance classifier $F$.
IV Proposed Model
@@
In this section, we describe our proposed model STANCE-C3. As illustrated in Figure 2, the architecture consists of two main components: (1) a domain counterfactual generator, which is trained with input representations of the labeled samples from the source domain and a small portion from the target domain, to generate a set of domain counterfactuals. These generated counterfactuals enrich the knowledge of the target domain during the stance classification training phase. (2) These samples are then combined with the original dataset (i.e., source domain dataset with a small portion of the target domain dataset) to train a BERT-based stance classifier. This training utilizes contrastive learning to develop a domain-adaptive, cross-target stance detection classifier. We partition a small subset (i.e., 50) of randomly selected samples from the combined datasets into positive and negative pairs based on their stance, regardless of the target. This contrastive loss approach enhances the method’s ability to generalize stance detection to unseen targets. The following subsections provide a detailed explanation of these two stages.
IV-A Domain-adaptive Single-target
@@
One major challenge in domain-adaptive stance detection is that collecting samples for a target domain can be expensive and usually requires human annotations. Even with annotated data, modeling a domain-adaptive stance detection is difficult due to the complexity of the dataset which contains far more distinct features over domains [37]. These features often form spurious correlations within a specific domain and make the models brittle to domain shift. For instance, the presence of question marks in a specific stance or highly negative sentiment associated with a particular stance has been demonstrated to lead to a performance decrease [6].
@@
To address the aforementioned challenges and control these variants on the domain-adaptive stance detection problem, we adapt counterfactual concepts to the source domain. A domain-counterfactual example is defined as a result text of intervening on the domain variable of the original example and converting it to the target domain while keeping other features (e.g., overall structure of the sentence) fixed. Given an example data from the COVID-19 vaccination domain (source domain), the generator first recognizes the terms related to the domain. Then, it intervenes on these terms, replacing them with text that links the example to the wearing face mask domain (target domain) while keeping the stance. We utilize source domain instances in combination with a small portion of target domain instances to incorporate target domain information in generating the counterfactuals specific to the target domain. These counterfactuals are then combined as input for the subsequent stage. Following the work of  [38], we adopt a T5-based language model to generate counterfactuals for a given target domain $\mathcal{T}$. In this approach, we train the T5-based LM to convert samples from source domain $\mathcal{S}$ to a given target domain $\mathcal{T}$.
@@
Generating domain counterfactuals consists of two main steps - domain corruption and reconstruction. For domain corruption, we get the masking score of all n-grams $w$ ($n\in\{1,2,3\}$) in a dataset $\mathcal{D}_{\mathcal{S}}$ following [38]. The affinity of $w$ to domain $\mathcal{S}$ is defined as $\rho(w,\mathcal{S})=P(\mathcal{S}|w)(1-\frac{H(\mathcal{S}|w)}{\log N})$, where $H(\mathcal{S}|w)$ is the entropy of $\mathcal{S}$ given each n-gram $w$ and $N$ is the number of unlabeled domains we know. The final masking score given source domain $\mathcal{S}$ and target domain $\mathcal{T}$ is $mask(w,\mathcal{S},\mathcal{T})=\rho(w,\mathcal{S})-\rho(w,\mathcal{T})$. A higher score implies the n-gram is highly related to the source domain and distant from the target domain.
@@
The next step involves reconstructing the masked examples $M(x)$ by concatenating them with a trainable embedding matrix, or domain orientation vectors. These vectors are embedded with domain-specific information including the nuances, vocabulary, and style of a particular domain. They are initialized with the embedding vector of each domain name and representative words from that domain. The concatenated matrix is then trained using a T5 architecture. Note that $\mathcal{T}=\mathcal{S}$ and $mask(w,\mathcal{S},\mathcal{T})=0$ during the training process. Given the target domain and its orientation vectors in the test phase, the trained model generates domain counterfactual $x^{\prime}$.
IV-B Domain-adaptive Cross-target